An airplane travels 600 miles against the wind in 4 hours, and makes the return trip with the same wind in 2 hours. Find the rate of the wind.
Answer by checkley77(12569)

SORRY I HASTILY DIVIDED 600/4 & GOT 200 INSTEAD OF 150!!!!!!!!!
------------------------------------------------------------------
600/4=150 mph against the wind.
600/2=300 mph with the wind.
300-150)/2=150/2=75 mph is the wind speed.
600/(150+75)=600/(300-75)
600/225=600/225